A confined of gas at 10 has a pressure of 2.5 atm. What is its new pressure if the temperature is increased to 60?





1
Expert's answer
2021-11-09T10:24:46-0500

"\frac{P_1}{T_1}=
\frac{P_2}{T_2}"


"P_2=\frac{P_1×T_2}{T_1}"


"P_2=\frac{2.5×[273+60]}{[273+10]}=2.94atm"